OpenStudy (anonymous):

limx-->25 ... sqrtx-5/x-25

OpenStudy (unklerhaukus):

\[\lim_{x\rightarrow25}\quad \frac{\sqrt{x-5}}{x-25}\] \[{^{L'H}}=\lim_{x\rightarrow25}\quad x^{-1/2}\]\[=25^{-1/2}\]\[=\frac{1}{ 25^{1/2}}\]\[=1/5\]